# Nes2Net: A Lightweight Nested Architecture for Foundation Model Driven Speech Anti-spoofing

### v1

@@ V-D Studies on the ASVspoof 2021 dataset
TABLE VI: Performance in EER (%) on the ASVspoof 2021 LA and DF. The results are reported in the format of ‘best (mean)’. CKPT Avg. refers to the number of checkpoints averaged. ‘Re-imp’ denotes re-implementation conducted by us. ‘Algo4’ and ‘Algo5’ represent Rawboost series augmentations: ‘(1+2+3)’ and ‘(1+2)’ [54], respectively. Parameters that are underlined are calculated by us. ‘-’ represents unknown. ‘N/A’ indicates that the system does not use the average checkpoints method.
@@
2023 wav2vec 2.0 Conformer [66] 2,51k [^5] https://github.com/ErosRos/conformer-based-classifier-for-anti-spoofing - 1.38 2.27
2024 wav2vec 2.0 Ensembling [67] - - 2.32 (4.48) 5.60 (8.74)
@@
re-imp wav2vec 2.0 AASIST (algo4) 447k N/A 1.13 (1.36) 3.37 (4.09)
re-imp wav2vec 2.0 AASIST (algo5) 447k N/A 0.93 (1.40) 3.56 (5.07)
Ours wav2vec 2.0 Nes2Net-X 511k N/A 1.73 (1.95) 1.65 (1.91)
@@
※: with extra data augmentation.
The results in Table VI show that for the LA track, our model achieves the best performance of 1.66% EER, comparable to SOTA systems. For the DF track, which more closely reflects real-world scenarios as discussed earlier, the baseline system AASIST [15] achieves a 2.85% EER, remaining competitive with current SOTA systems. The SLS [35] and TCM [33] models achieve EERs close to 2%, demonstrating strong performance at the SOTA level. The Mamba-based [17] model further improves results, reducing the EER to 1.88%. Notably, our proposed Nes2Net-X achieves the best performance among all compared systems, with an EER of 1.65%.
TABLE VII: Performance in EER (%) for different types of vocoders and compression conditions on the ASVspoof 2021 DF test set. The five EER values for each sub-item, from left to right, correspond to Nes2Net-X, Mamba [17], SLS [35], TCM [33], and AASIST [15].The best performance is reported in bold fonts, and the second-best is underlined.
@@
Inspired by prior works [33, 17], we further experiment by averaging the weights of the top-performing validation checkpoints to produce an improved model. This approach further improves the performance of the DF task to an EER of 1.49%, which, to the best of our knowledge, is the best performance reported to date. Furthermore, compared to Mamba [17], our model achieves this performance with approximately 74% fewer parameters, demonstrating superior efficiency.
The analysis above summarizes overall performance on the DF test set. The DF dataset also provides detailed labels for vocoder types and compression conditions, enabling more fine-grained analysis. To further evaluate performance, we compare the SOTA models—Mamba, SLS, TCM, and AASIST—with our proposed Nes2Net-X across these sub-tracks. The results are presented in Table VII. To improve readability and make the extensive numerical data easier to interpret, we also visualize the table’s results in Fig. 4.
@@
From the perspective of compression conditions, the differences in model performance are less pronounced compared to the variations observed across vocoder types. Nes2Net-X consistently achieves the lowest EERs across all compression conditions, regardless of the level of distortion introduced by compression. This consistency highlights the model’s adaptability and strong generalization ability across different levels of compressions.
@@
V-E The results on the In-the-Wild dataset
TABLE VIII: Performance in EER (%) on the In-the-Wild [50] dataset. Our result is reported as the format of ‘best (mean)’ across 3 runs.
@@ V-E The results on the In-the-Wild dataset
wav2vec 2.0 2022 AASIST [15] 10.46
2024 SLIM [14] 12.5
@@
※: with extra data augmentation.
@@
In addition, unlike many other datasets that rely on self-generated fake audio, this dataset is collected from publicly available video and audio files explicitly labeled as audio deepfakes [50]. To account for the potential presence of partial spoofing, we evaluate our proposed Nes2Net-X using the entire duration of each test sample instead of restricting it to the first 4 seconds, as the latter approach risks missing partially spoofed segments.
The testing results, alongside SOTA models, are reported in Table VIII. We find that the overall performance trends are consistent with those seen on the ASVspoof 2021 DF dataset. However, EERs on the In-the-Wild dataset are generally higher than those on the DF dataset, reflecting greater complexity and variability in real-world scenarios. Notably, the proposed Nes2Net-X outperforms all SOTA models, achieving the lowest EER of 5.52% on this challenging dataset.
TABLE IX: The performance in EER (%) on the ASVspoof 2021 LA, DF [48], and In-the-Wild [50] datasets. The results are reported as the format of ‘best (mean)’ across 3 runs. ‘w/ Aug.’ and ‘w/o Aug.’ indicate whether evaluation with augmentations on the validation set is used to select the best checkpoint for testing. CKPT Avg. refers to the number of checkpoints averaged.
Back-end Train Set CKPT Avg. w/ Aug. w/o Aug.
21LA [48] 21DF [48] In-the-Wild [50] 21LA [48] 21DF [48] In-the-Wild [50]
3 1.70 (1.80) 1.88 (1.98) 5.15 (6.31) 1.66 (1.87) 1.54 (1.98) 5.59 (6.90)
5 1.67 (1.78) 1.80 (1.91) 5.28 (6.31) 1.88 (2.00) 1.49 (1.78) 5.52 (6.60)
V-F The results on the ASVspoof 5 dataset
TABLE X: A comparison between the proposed Nes2Net and the AASIST baseline system on the ASVspoof 5 dataset [49]. ‘Params.’ and ‘MMACs’ refer to the number of parameters and the number of million multiply-accumulate operations, respectively. ‘Avg.’ indicates the average relative performance improvement across all three evaluation metrics.
@@ V-F The results on the ASVspoof 5 dataset
The ASVspoof 5 dataset represents the most recent edition in the ASVspoof series. Unlike earlier versions, it introduces adversarial attacks and is crowdsourced under various acoustic conditions [49]. As it is newly released, there are currently no existing systems available for a fair comparison. Therefore, using the WavLM front-end, we implement the AASIST system as a baseline and compare it with our proposed Nes2Net-X model. Based the evaluation protocol in  [37], we assess performance using three metrics: Cost of Log-Likelihood Ratio (CLLR), minimum Detection Cost Function (minDCF), and EER, and present the results in Table X. We observe that the Nes2Net-X back-end model results in only a slight increase in the number of parameters compared to AASIST, while significantly reducing MMACs. Moreover, across all three evaluation metrics, the Nes2Net-X back-end improves performance by 10.9%.
V-G The results on the PartialSpoof dataset
TABLE XI: Performance in EER(%) on the PartialSpoof [51] dataset. The results are reported as the format of ‘best (mean)’ across 3 runs. † indicates results obtained from our implementation.
@@ V-G The results on the PartialSpoof dataset
Partially manipulating a sentence can significantly alter its intended meaning [58]. When such manipulations occur in small regions, existing models—trained on fully spoofed speech and relying on pooling functions—struggle to detect these subtle changes. Consequently, there is growing interest in the detection of partially spoofed speech [58, 51, 78].
To evaluate the performance of our proposed model across different datasets and spoofing tasks, we conduct experiments on the PartialSpoof dataset [51]. The results are presented in Table XI. First, we reproduce the performance of two SOTA models, achieving results comparable to those reported in their original papers [58, 51]. Additionally, we evaluate SE ResNet, which demonstrated performance similar to the other baselines. In contrast, our proposed Nes2Net and Nes2Net-X significantly outperform all three baselines, achieving relative performance improvements of 36.6% and 25.6%, respectively.
@@ V-H Should We Use Augmentation During Validation?
In this section, we compare these two approaches and report the results in Table IX. We observe that applying the same augmentation to the validation set as in the training set leads to worse performance on ASVspoof 2021 DF [48], but better results on In-the-Wild [50]. When no augmentation is applied to the validation set, the opposite trend is observed.
@@ VI Conclusion
In this work, we propose Nested Res2Net (Nes2Net) and its enhanced variant, Nes2Net-X, as lightweight and dimensionality reduction (DR) layer-free back-end architectures designed for speech anti-spoofing in the foundation model era. Unlike conventional approaches that rely on a DR layer to bridge the mismatch between high-dimensional features and downstream classifiers, our proposed architectures directly process these rich representations. This not only eliminates the computational and parameter burden introduced by DR layers but also avoids information loss, enhancing overall system efficiency and robustness.
@@
We conduct extensive evaluations across five representative datasets—CtrSVDD, ASVspoof 2021, ASVspoof 5, PartialSpoof, and In-the-Wild—covering a wide range of singing voice deepfakes, fully spoofed speech, adversarial attack, real-world deepfakes, and partially spoofed speech. Across all scenarios, our models achieve SOTA performance, demonstrating superior generalization, compactness, and resilience under unseen and challenging conditions.
In summary, Nes2Net and Nes2Net-X offer a general-purpose, resource-efficient back-end for foundation model-based speech anti-spoofing, providing a practical yet powerful alternative to DR-dependent designs. To support future research and application, we make all source code and pre-trained models publicly available.

### v2

@@ V-C Studies on the ASVspoof 2021 dataset
TABLE V: Performance in EER (%) on the ASVspoof 2021 LA and DF. The results are reported in the format of ‘best (mean)’. CKPT Avg. refers to the number of checkpoints averaged. $\dagger$}’ denotes re-implementation conducted by us. ‘Algo4’ and ‘Algo5’ represent Rawboost series augmentations: ‘(1+2+3)’ and ‘(1+2)’ [55], respectively. Parameters that are underlined are calculated by us. ‘-’ represents unknown. ‘N/A’ indicates that the system does not use the average checkpoints method.
@@
2023 wav2vec 2.0 Conformer (fix) [63] 2,506k [^5] https://github.com/ErosRos/conformer-based-classifier-for-anti-spoofing 5 1.38 2.27
2023 wav2vec 2.0 Conformer (var) [63] 2,506k 5 0.87 7.36
2024 wav2vec 2.0 Ensembling [64] $\ddagger$ - - 2.32 (4.48) 5.60 (8.74)
@@
$\dagger$} wav2vec 2.0 AASIST (algo4) 447k N/A 1.13 (1.36) 3.37 (4.09)
$\dagger$} wav2vec 2.0 AASIST (algo5) 447k N/A 0.93 (1.40) 3.56 (5.07)
Ours wav2vec 2.0 Nes2Net 511k N/A 1.61 (1.90) 1.89 (2.12)
wav2vec 2.0 Nes2Net-X 511k N/A 1.73 (1.95) 1.65 (1.91)
@@
※: with extra data augmentation [71]
$\ddagger$: ensemble of multiple models
TABLE VI: Performance in EER (%) for different types of vocoders and compression conditions on the ASVspoof 2021 DF test set. The five EER values for each sub-item, from left to right, correspond to Nes2Net-X, Mamba [17], SLS [35], TCM [33], and AASIST [15].The best performance is reported in bold fonts, and the second-best is underlined.
@@
The ASVspoof series datasets are widely used as benchmarks for advancing research in detecting spoofed speech [47, 48]. Following the standard protocol, we train models on ASVspoof 2019 [47] and evaluate them on ASVspoof 2021 Logical Access (LA) and Deepfake (DF) tasks [48]. The LA task focuses on detecting synthetic and voice-converted speech transmitted over telephony systems, introducing challenges related to channel effects and transmission variability. In contrast, the DF task targets detecting manipulated, compressed speech data commonly found on online platforms. This reflects real-world scenarios where deepfake audio circulates, making the DF task a valuable benchmark for evaluating deepfake detection systems.
The results in Table V show that for the LA track, our Nes2Net achieves a mean EER of 1.90%, comparable to SOTA systems. For the DF track, which more closely reflects real-world scenarios as discussed earlier, the baseline system AASIST [15] achieves its best EER of 2.85% and a mean EER of 3.69%, remaining competitive with current SOTA systems. The SLS [35] and TCM [33] models achieve EERs close to 2%, demonstrating strong performance at the SOTA level. The Mamba-based [17] model further improves results, reducing the EER to 1.88%. Notably, our proposed Nes2Net attains its best EER of 1.89% and a mean EER of 2.12% EER, comparable to the performance of current SOTA systems. The enhanced variant, Nes2Net-X achieves the best performance among all compared systems, with its best EER of 1.65% and a mean EER of 1.91%.
Inspired by prior works [33, 17], we average the weights of several top-performing checkpoints on the validation set to obtain an improved model. This approach further improves the performance of the DF task to a best EER of 1.49% and a mean EER of 1.78%, which, to the best of our knowledge, is the best performance reported to date. Furthermore, compared to Mamba [17], our model achieves this performance with approximately 74% fewer parameters, demonstrating superior efficiency.
The analysis above summarizes overall performance on the DF test set. The DF dataset also provides detailed labels for vocoder types and compression conditions, enabling more fine-grained analysis. To further evaluate performance, we compare the SOTA models Mamba, SLS, TCM, and AASIST with our proposed Nes2Net-X across these sub-tracks. The results are presented in Table VI. To improve readability and make the extensive numerical data easier to interpret, we also visualize the table’s results in Fig. 4.
@@
From the perspective of compression conditions, the differences in model performance are less pronounced compared to the variations observed across vocoder types. Nes2Net-X consistently achieves the lowest EERs across all compression conditions, regardless of the level of distortion introduced by compression. This consistency highlights the model’s strong generalization ability across different levels of compressions.
@@
V-D The results on the In-the-Wild dataset
TABLE VII: Performance in EER (%) on the In-the-Wild [50] dataset. Our result is reported as the format of ‘best (mean)’ across 3 runs.
@@ V-D The results on the In-the-Wild dataset
wav2vec 2.0
2022 Rawnet&ASSIST (reported by [35]) 10.46
2024 SLIM [14]    - (12.5)
@@
※: with extra data augmentation [71]
@@
In addition, unlike many other datasets that rely on self-generated fake audio, this dataset is collected from publicly available video and audio files explicitly labeled as audio deepfakes [50]. To account for the potential presence of partial spoofing, we evaluate our proposed Nes2Net and Nes2Net-X using the entire duration of each test sample instead of restricting it to the first 4 seconds, as the latter approach risks missing partially spoofed segments.
The testing results, alongside SOTA models, are reported in Table VII. We find that the overall performance trends are consistent with those seen on the ASVspoof 2021 DF dataset. However, EERs on the In-the-Wild dataset are generally higher than those on the DF dataset, reflecting greater complexity and variability in real-world scenarios. Notably, the proposed Nes2Net-X outperforms all SOTA models, achieving the lowest EER of 5.52% and a mean EER of 6.60% on this challenging dataset.
V-E The results on the ASVspoof 5 dataset
TABLE VIII: A comparison between the proposed Nes2Net and the AASIST baseline system on the ASVspoof 5 dataset [49]. ‘Params.’ and ‘MMACs’ refer to the number of parameters and the number of million multiply-accumulate operations, respectively. ‘Avg.’ indicates the average relative performance improvement across all three evaluation metrics.
@@ V-E The results on the ASVspoof 5 dataset
The ASVspoof 5 dataset represents the most recent edition in the ASVspoof series. Unlike earlier versions, it introduces adversarial attacks and is crowdsourced under various acoustic conditions [49]. As it is newly released, there are currently no existing systems available for a fair comparison. Therefore, we re-implement the AASIST system as a baseline and compare it with our proposed Nes2Net and Nes2Net-X model. Following the ASVspoof 5 challenge guidelines [49], we use WavLM [3] as the front-end. Based the evaluation protocol in [37], we assess performance using three metrics: Cost of Log-Likelihood Ratio (CLLR), minimum Detection Cost Function (minDCF), and EER, and present the results in Table VIII. We observe that the Nes2Net and Nes2Net-X back-end models result in only a slight increase in the number of parameters compared to AASIST, while significantly reducing MMACs. Moreover, across all three evaluation metrics, the Nes2Net and Nes2Net-X back-ends improve performance by 7.1% and 10.9%, receptively.
V-F The results on the PartialSpoof dataset
TABLE IX: Performance in EER(%) on the PartialSpoof [51] dataset. The results are reported as the format of ‘best (mean)’ across 3 runs. † indicates results obtained from our implementation.
@@ V-F The results on the PartialSpoof dataset
Partially manipulating a sentence can significantly alter its intended meaning [59]. When such manipulations occur in small regions, existing models trained on fully spoofed speech and relying on pooling functions struggle to detect these subtle changes. Consequently, there is growing interest in the detection of partially spoofed speech [59, 51, 75].
To evaluate the performance of our proposed model across different spoofing tasks, we conduct experiments on the PartialSpoof dataset [51]. The results are presented in Table IX. First, we reproduce the performance of two SOTA models, achieving results comparable to those reported in their original papers [59, 51]. Additionally, we evaluate SE ResNet, which demonstrated performance similar to the other baselines. In contrast, our proposed Nes2Net and Nes2Net-X outperform all three baselines.
TABLE X: The performance in EER (%) on the ASVspoof 2021 LA, DF [48], and In-the-Wild [50] datasets. The results are reported as the format of ‘best (mean)’ across 3 runs. ‘w/ Aug.’ and ‘w/o Aug.’ indicate whether evaluation with augmentations on the validation set is used to select the best checkpoint for testing. CKPT Avg. refers to the number of checkpoints averaged.
Back-end Train Set
CKPT
Avg.
w/ Aug. w/o Aug.
21LA [48] 21DF [48] In-the-Wild [50] 21LA [48] 21DF [48] In-the-Wild [50]
3 1.70 (1.80) 1.88 (1.98) 5.15 (6.31) 1.66 (1.87) 1.54 (1.98) 5.59 (6.90)
5 1.67 (1.78) 1.80 (1.91) 5.28 (6.31) 1.88 (2.00) 1.49 (1.78) 5.52 (6.60)
V-G Empirical Runtime and Memory Analysis
Number of parameters and MMACs are widely adopted metrics for evaluating model efficiency. These platform-independent measures offer consistent and fair comparisons across different hardware. However, to better reflect the real-world deployment costs of back-end architectures, we additionally benchmark their training time, inference time, and peak GPU memory usage, as summarized in Table XI.
TABLE XI: Training and inference efficiency comparison across back-end models. The table reports the average (Avg.) training and inference time per batch in milliseconds (ms/batch), as well as peak GPU memory usage in megabytes (MB).
Back-end Avg. Time (ms/batch)↓ Peak GPU Memory↓
Training Inference (MB)
AASIST Light (C=24) 27.0 7.8 1,327
AASIST Standard(C=32) 53.8 18.7 3,454
AASIST Large(C=40) 79.2 28.1 4,273
AASIST XL(C=48) 86.1 30.7 5,087
AASIST XXL(C=56) 100.9 37.4 5,905
ResNet 7.8 2.6 691
Res2Net 15.6 3.5 721
ECAPA-TDNN (C=128) 9.4 3.1 698
Proposed Nes2Net 20.2 4.9 1,312
Proposed Nes2Net-X 29.1 9.2 2,231
All back-end models are evaluated under identical conditions: input features of 400 frames with 1024 dimensions, a batch size of 64, and execution on a dedicated NVIDIA H20 GPU. The first 10 batches are used for warm-up and excluded from the measurement, and the inference and training times are averaged over the subsequent 200 batches. Training time includes the forward, backward, and optimizer update steps.
The results show that AASIST models exhibit rapidly increasing runtime and memory consumption as the channel dimension $C$ grows. In contrast, our proposed Nes2Net achieves notably lower latency and memory usage. Nes2Net-X further improves performance in some settings by preserving more high-dimensional information, albeit at the cost of higher resource consumption.
Conventional models such as ResNet, Res2Net, and ECAPA-TDNN offer faster runtime and smaller memory footprints than our proposed method, but fall short in detection accuracy as shown in earlier experiments. Therefore, when selecting a back-end architecture, we believe both Nes2Net and Nes2Net-X offer flexible options: the former prioritizes efficiency, while the latter favors accuracy when computational resources permit. This underscores the importance of balancing performance and efficiency in real-world applications.
@@ V-H Should We Use Augmentation During Validation?
In this section, we compare these two approaches and report the results in Table X. We observe that applying the same augmentation to the validation set as in the training set leads to worse performance on ASVspoof 2021 DF [48], but better results on In-the-Wild [50]. When no augmentation is applied to the validation set, the opposite trend is observed.
@@ VI Conclusion
In this work, we propose Nested Res2Net (Nes2Net) and its enhanced variant, Nes2Net-X, as lightweight and dimensionality reduction (DR) layer-free back-end architectures designed for speech anti-spoofing in the era of foundation models. Unlike conventional approaches that rely on a DR layer to bridge the mismatch between high-dimensional features and downstream classifiers, our proposed architectures directly process these rich representations. This not only eliminates the computational and parameter overhead introduced by DR layers but also avoids information loss, enhancing overall system efficiency and robustness.
@@
We conduct extensive evaluations across five representative datasets: CtrSVDD, ASVspoof 2021, ASVspoof 5, PartialSpoof, and In-the-Wild, covering a wide range of singing voice deepfakes, fully spoofed speech, adversarial attacks, real-world deepfakes, and partially spoofed speech. Across all scenarios, our models achieve SOTA performance, demonstrating superior generalization, compactness, and resilience under unseen and challenging conditions.
In summary, Nes2Net and Nes2Net-X offer a general-purpose, resource-efficient back-end for foundation model-based speech anti-spoofing, providing a practical yet powerful alternative to DR-dependent designs. To facilitate future research and applications, we make all source code and pre-trained models publicly available.